### v3

@@ 3.1 Task Formulation
Financial Concept Linking (FinCL)
@@
$f_{\text{FinCL}}:(e,l,C_{e},\mathcal{T})\mapsto\hat{c}$ (3)
@@
3.2 Raw Data Collection
We compiled 142 annual 10-K filings from publicly listed U.S. companies submitted in 2023–2024 (Table 2). To ensure coverage diversity, the collection spans all 11 major industry sectors, with about 95% of companies located across 31 U.S. states and the remainder based internationally. Using BeautifulSoup, we parsed the reports and extracted 319,893 narrative sentences and 21,576 financial tables. Company-level statistics are provided in Appendix C.1. All filings follow the SEC Inline XBRL (iXBRL) format, which embeds machine-readable tags inside HTML disclosures and links textual and tabular values to US-GAAP taxonomy concepts through schema and linkbase definitions. However, these regulatory tags are designed for compliance rather than model evaluation. They provide little coverage of negative cases and contain substantial structural noise. We therefore construct a high-fidelity evaluation suite that repurposes this raw regulatory metadata into a standardized benchmark for financial fact extraction and concept linking. Additional discussion of our benchmark granularity is provided in Appendix C.2.
Table 2: Statistics of collected financial reports.
@@ 3.2 Raw Data Collection
3.3 Large-Scale Automated Annotation
@@
Table 3: Statistical information for the original dataset in our benchmark. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
@@
Automated Annotation Engine. We build a rule-based parser (Algorithm 1, Appendix C.3) that reconstructs semantic links between human-readable disclosures and Inline XBRL (iXBRL) metadata. For each instance $i$, the parser locates ix:nonfraction elements and extracts the value $v$, concept $c$, and data type $\tau$, while restricting extraction to atomic leaf elements $\epsilon$ to avoid duplicates from nested tags.
@@
Structural Refinement and Filtering. For reliable evaluation, we keep only text segments longer than 20 characters and deduplicate content globally using $\mathcal{S}_{text}$. For tables, we retain structural tags (<table>, <tr>, <td>, <th>) but discard tables without numerical facts. This filtering preserves semantic accuracy while maintaining realistic layout structure.
Figure 2: Statistics of numerical entity types.
3.4 Data Quality Control
Validating all 261,457 entities is infeasible. To obtain a rigorous quality upper bound, we adopt a strategic sampling scheme based on a greedy minimal-cover algorithm (Algorithm 2, Appendix D.1). Rather than random sampling, we construct a diversity-maximized subset $\mathcal{P}^{*}$ that satisfies two goals: (1) Company Balance, limiting bias from individual reporting styles, and (2) Taxonomy Completeness, ensuring that each US-GAAP concept observed in the corpus appears at least once. We seed $\mathcal{P}^{*}$ with capped samples per company, then iteratively add instances that cover the largest number of uncovered concepts, producing a compact but representative audit set. The final subset contains 928 positive sentence instances and 1,021 positive table instances, serving as a taxonomy-wide stress test reviewed by two junior analysts and one senior adjudicator (the demography of these annotators is shown in Appendix D.2).
Review Protocol. The audit followed two stages. A pilot phase aligned interpretations of the Verification Guideline (Appendix D.3), after which analysts independently verified whether extracted triplets $(e,\tau,c)$ were fully correct. The senior expert adjudicated disagreements to create the final gold set.
Agreement Results. On the independent round, annotators achieved 96% raw agreement and Cohen’s $\kappa=0.81$, indicating substantial agreement landis1977measurement, the equation is shown in Appendix D.4. This high-verified accuracy on the most semantically diverse subset provides strong evidence for the reliability of our automated annotation pipeline.
3.5 Task-Specific Dataset Construction
@@
Table 4: Statistics of evaluation datasets for FinNI-eval and FinCL-eval. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
@@
3.5.1 FinNI-eval Dataset Construction
Based on the annotation procedure described in Section 3.3, we construct the FinNI-eval dataset, comprising 28,787 instances derived from the annotated sentences and tables. Specifically, the dataset comprises two components: the input block and the answer block. The input block includes a task instruction and the input content. The instruction defines the FinNI task by specifying the responsibilities of the LLMs, the definitions of entity types, and the rules to be followed. The answer block is formatted as a list aligned with the input block. The answer is a JSON list of all identified entity values and types or an empty JSON list. The specific prompt template is shown in Figure 9.
@@ 3.5.1 FinNI-eval Dataset Construction
3.5.2 FinCL-eval Dataset Construction
Following the formulation of the FinCL subtask in Section 3.1, we further construct a FinCL-eval dataset that includes 261,457 query-answer pairs for numerical entity normalization. In addition, we built a US-GAAP taxonomy database containing 17,688 unique financial concepts. To create the FinCL-eval dataset, we utilize all positive instances identified in Section 3.3. Each query consists of a numerical entity, its corresponding entity type, and its surrounding context, while the answer is the associated US-GAAP concept.
@@ 3.6.1 Evaluation Framework
We propose a unified evaluation framework for 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:. As illustrated in Figure 3, each input is decomposed into two subtasks, FinNI and FinCL, which together produce structured triplets $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$. This design jointly evaluates zero-shot fact extraction and concept alignment. For comparison with token-classification baselines, we report macro and micro Precision, Recall, and F1 sokolova2009systematic. FinNI is evaluated using pair-level metrics, and FinCL with accuracy. Full metric definitions are in Appendix E.
Figure 3: A unified evaluation framework on 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark. Note: The Input content is drawn from the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: dataset, and the “Shared LLMs” are used for identification in the FinNI task and reranking in the FinCL task.
FinNI Evaluation. We adopt pair-level metrics to measure extraction quality, focusing on whether models can locate financial values from mixed text–table contexts based on semantic cues.
FinCL Evaluation. To test fine-grained numeric entity normalization, we formulate FinCL as a retrieval–reranking task: a retrieval stage produces a candidate set of taxonomy concepts, and the LLM then reranks candidates to select the most appropriate tag, avoiding impractical extreme classification (prompt template in Figure 10, Appendix).
@@
To this end, we embed each taxonomy concept using text-embedding-3-small [^5] https://platform.openai.com/docs/models/text-embedding-3-small , then retrieve the top-$h$ candidate tags from the US-GAAP taxonomy $\mathcal{T}$ by semantic similarity. The LLM subsequently reranks these candidates using contextual reasoning to select the best tag $\hat{c}\in\mathcal{T}_{h}$:
$f_{\text{rerank}}:(e,l,C_{e},\mathcal{T}_{h})\mapsto\hat{c}$ (4)
3.6.2 Evaluation Models
Our goal is to assess the foundational capabilities of SOTA LLMs on 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: and understand their strengths and limitations in XBRL tagging. We evaluate models across three categories using our unified framework (Table 11). (1) Closed-source general LLM: GPT-4o hurst2024gpt. (2) Open-source general LLMs: DeepSeek-V3 liu2024deepseek, DeepSeek-R1-Distill-Qwen-32B deepseekai2025deepseekr1incentivizingreasoningcapability, Qwen3-32B / 14B / 1.7B / 0.6B qwen3technicalreport, Llama-4-Scout-17B-16E-Instruct meta2025llama, Llama-3.3-70B-Instruct, Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct grattafiori2024llama, and gemma-2-27b-it gemma_2024. (3) Domain-specific financial LLM: Fino1-8B qian2025fino1. We also compare against strong PLM baselines: BERT-large DBLP:journals/corr/abs-1810-04805, FinBERT araci2019finbert, and SECBERT loukas2022finer.
3.6.3 Evaluation setting
@@
4 Experiment and Result
4.1 Overall Results
@@
Table 5: Overall Performance. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
Category Models Macro Micro
@@
Closed-source LLM GPT-4o 0.0865 0.0626 0.0569 0.1052 0.0731 0.0863
Open-source LLMs DeepSeek-V3 0.0949 0.0778 0.0659 0.1074 0.1264 0.1162
@@
Financial LLM Fino1-8B 0.0344 0.0143 0.0151 0.0419 0.0128 0.0197
Fine-tuned PLMs BERT-large 0.0252 0.0266 0.0205 0.1518 0.1283 0.1391
@@
From a macro perspective, which emphasizes balanced performance across frequent and rare tags, large general-purpose LLMs clearly dominate. DeepSeek-V3, GPT-4o, and Llama-4-Scout achieve the strongest macro-F1 scores, surpassing all fine-tuned PLMs and indicating better generalization to long-tail concepts. Qwen3-32B also performs competitively, suggesting that model architecture and pretraining can compensate for a smaller scale. From a micro perspective, which reflects performance on frequent labels, the trend is similar: top LLMs remain competitive even without fine-tuning, while fine-tuned PLMs such as SECBERT and BERT-large excel mainly because they benefit from label frequency rather than broader coverage (Table 5).
4.2 Subtask Results
4.2.1 FinNI subtask
Table 6 summarizes results on FinNI. Overall, larger LLMs show markedly stronger ability to identify numerical entities and produce structured outputs, even without financial adaptation. DeepSeek-V3 achieves the best overall performance, while GPT-4o performs competitively and ranks among the top models.
@@ 4.2.1 FinNI subtask
Table 6: Performance comparison of different models on the FinNI subtask. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
Category Model Precision Recall F1
Closed-source LLM GPT-4o 0.5893 0.4977 0.5397
Open-source LLMs DeepSeek-V3 0.5886 0.8430 0.6932
Llama-4-Scout-17B-16E-Instruct 0.4668 0.3164 0.3771
Llama-3.3-70B-Instruct 0.4826 0.3301 0.3920
DeepSeek-R1-Distill-Qwen-32B 0.5676 0.1942 0.2894
Qwen3-32B 0.6991 0.1804 0.2868
gemma-2-27b-it 0.5060 0.4526 0.4778
Qwen3-14B 0.6912 0.1487 0.2448
Llama-3.1-8B-Instruct 0.3874 0.1761 0.2421
Llama-3.2-3B-Instruct 0.1856 0.1203 0.1460
Qwen3-1.7B 0.7362 0.0281 0.0541
Qwen3-0.6B 0.2803 0.0019 0.0038
Financial LLM Fino1-8B 0.3431 0.1293 0.1878
A notable pattern is that some models (e.g., Qwen3 variants) obtain very high precision but extremely low recall, indicating that they extract only a small fraction of target facts. This confirms that precision alone is insufficient for FinNI, where missing facts are often more damaging than producing a few spurious ones.
Smaller open-source models struggle severely, reflecting the difficulty of this zero-shot extraction setting. In contrast, mid-size models such as gemma-2-27B achieve more balanced behavior. Interestingly, the financial LLM Fino1-8B still lags behind general-purpose models, suggesting that domain text pretraining alone does not help unless the training task aligns closely with FinNI.
4.2.2 FinCL subtask
Table 7 reports FinCL performance under the reranking setup, using candidates retrieved from the best top-200 configuration (Table 14). Although retrieval provides essential coverage, low Acc@k values, especially for table entities, substantially cap downstream accuracy.
Table 7: Performance comparison of different models on the FinCL subtask. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
Category Model Accuracy
Closed-source LLM GPT-4o 0.1829
Open-source LLMs DeepSeek-V3 0.1889
Llama-4-Scout-17B-16E-Instruct 0.1649
Llama-3.3-70B-Instruct 0.1318
DeepSeek-R1-Distill-Qwen-32B 0.1141
Qwen3-32B 0.1277
gemma-2-27b-it 0.1099
Qwen3-14B 0.1144
Llama-3.1-8B-Instruct 0.0913
Llama-3.2-3B-Instruct 0.0415
Qwen3-1.7B 0.0735
Qwen3-0.6B 0.0414
Financial LLM Fino1-8B 0.0807
Within this setting, large LLMs such as DeepSeek-V3, GPT-4o, and Llama-4-Scout perform best, but overall accuracy remains low, reflecting the difficulty of aligning evidence with fine-grained taxonomy concepts. Most other models cluster well below this tier, and smaller models approach random behavior. The financial model Fino1-8B also lags behind general LLMs, suggesting that domain pretraining alone is insufficient without task-aligned supervision.
Taken together, FinCL exposes two compounding challenges: taxonomy disambiguation requires reasoning over subtle semantic cues, and retrieval errors propagate directly into reranking, creating a ceiling effect on end-to-end alignment.
4.3 Ablation analysis

### v4

@@ 3.1.1. Financial Numeric Identification (FinNI)
3.1.2. Financial Concept Linking (FinCL)
@@
(3) $f_{\text{FinCL}}:(e,l,C_{e},\mathcal{T})\mapsto\hat{c}$
@@
3.2. Raw Data Collection
We compiled 142 annual 10-K filings from publicly listed U.S. companies submitted to the U.S. Securities and Exchange Commission (SEC) [^5] https://www.sec.gov/ in 2023–2024 (Table 2). The collection covers all 11 major industry sectors, with companies primarily based in the United States and a small portion internationally. Parsing the filings with BeautifulSoup yields 319,893 narrative sentences and 21,576 financial tables. All filings follow the SEC Inline XBRL (iXBRL) format, which embeds machine-readable tags within HTML disclosures and links textual and tabular values to US-GAAP taxonomy concepts. However, these regulatory tags are designed for compliance rather than model evaluation, providing limited negative coverage and substantial structural noise. We therefore repurpose the raw iXBRL metadata to construct a standardized, high-fidelity evaluation suite for financial fact extraction and concept linking.
Table 2. Statistics of collected financial reports.
@@ 3.2. Raw Data Collection
3.3. Large-Scale Automated Annotation
@@
Table 3. Statistical information for the original dataset in our benchmark. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
@@
Automated Annotation Engine. We build a rule-based parser that reconstructs semantic links between human-readable disclosures and Inline XBRL (iXBRL) metadata. For each instance $i$, the parser extracts iXBRL fact elements (e.g., ix:nonfraction) and retrieves the value $v$, concept $c$, and data type $\tau$, implemented as ParseIXBRL in Algorithm 3. To avoid duplicates from nested tags, we restrict extraction to atomic leaf elements $\epsilon$.
Input: Raw iXBRL reports $\mathcal{D}$; taxonomy $\mathcal{T}$; valid types $\mathcal{L}$
Initialize $\mathcal{P}\!\leftarrow\!\emptyset$, $\mathcal{N}\!\leftarrow\!\emptyset$, seen text set $\mathcal{S}\!\leftarrow\!\emptyset$;
foreach instance $i\in\mathcal{D}$ (sentence/table) do
  $text\leftarrow\textsc{ExtractText}(i)$;
  if $|text|\leq 20$ or $text\in\mathcal{S}$ then
    continue
 $E\leftarrow\{(v,\tau,c)\mid(v,c,\tau)\!\leftarrow\!\textsc{ParseIXBRL}(i),\ c\!\in\!\mathcal{T},\ \tau\!\in\!\mathcal{L},\ v\neq\emptyset\}$;
  $E\leftarrow\textsc{DedupBy}(E,\ (v,c))$;
 if $E\neq\emptyset$ then
     $\mathcal{P}\leftarrow\mathcal{P}\cup\{(text,E)\}$;
 else
     $\mathcal{N}\leftarrow\mathcal{N}\cup\{(text,\emptyset)\}$;
 $\mathcal{S}\leftarrow\mathcal{S}\cup\{text\}$;
return $\mathcal{P},\mathcal{N}$;
Algorithm 1 Automated Instance Annotation
@@
Structural Refinement and Filtering. For reliable evaluation, we keep only text segments longer than 20 characters and deduplicate content globally using texttext. For tables, we retain structural tags (<table>, <tr>, <td>, <th>) but discard tables without numerical facts. This filtering preserves semantic accuracy while maintaining realistic layout structure.
Figure 2. Statistics of numerical entity types.
3.4. Data Quality Control
Validating all 261,457 entities is infeasible. We therefore adopt a strategic sampling scheme based on a greedy minimal-cover method (algorithm 4) to construct a compact audit set $\mathcal{P}^{*}$. The procedure balances two objectives: Company Balance, by capping the number of sampled instances per company, and Taxonomy Completeness, by ensuring coverage of all observed US-GAAP concepts. Specifically, we first seed $\mathcal{P}^{*}$ with up to $K$ instances per company, then iteratively add instances that cover the largest number of remaining uncovered concepts. This process yields a diversity-maximized subset consisting of 928 sentence instances and 1,021 table instances, which is reviewed by two junior analysts and one senior adjudicator.
Input: Positive instances $\mathcal{P}$; observed concepts $\mathcal{C}_{all}$; cap $K=10$
Output: Strategic subset $\mathcal{P}^{*}$
$\mathcal{P}^{*}\!\leftarrow\!\emptyset$; $\mathcal{C}_{rem}\!\leftarrow\!\mathcal{C}_{all}$;
foreach company $g$ do
  Add up to $K$ instances from $g$ to $\mathcal{P}^{*}$;
  $\mathcal{C}_{rem}\leftarrow\mathcal{C}_{rem}\setminus\textsc{Concepts}(\mathcal{P}^{*})$;
while $\mathcal{C}_{rem}\neq\emptyset$ and $\mathcal{P}\setminus\mathcal{P}^{*}\neq\emptyset$ do
  $i^{*}\leftarrow\arg\max_{i\in\mathcal{P}\setminus\mathcal{P}^{*}}|\textsc{Concepts}(i)\cap\mathcal{C}_{rem}|$;
  if ties then
    select $i^{*}$ with the shortest text length;
 $\mathcal{P}^{*}\leftarrow\mathcal{P}^{*}\cup\{i^{*}\}$; $\mathcal{C}_{rem}\leftarrow\mathcal{C}_{rem}\setminus\textsc{Concepts}(i^{*})$;
return $\mathcal{P}^{*}$;
Algorithm 2 Strategic Audit Set Sampling
Review Protocol. The audit followed two stages. A pilot phase aligned interpretations of the Verification Guideline [^6] https://docs.google.com/document/d/1TgTxz-fozBImeRs8Y9kLViHnr8Bjs1qZbsB8u5NjfmA/edit?usp=sharing, after which analysts independently verified whether extracted triplets $(v,\tau,c)$ were fully correct. The senior expert adjudicated disagreements to create the final gold set.
Agreement Results. On the independent round, annotators achieved 96% raw agreement and Cohen’s $\kappa=0.81$, indicating substantial agreement (Landis and Koch, 1977). This high-verified accuracy on the most semantically diverse subset provides strong evidence for the reliability of our automated annotation pipeline.
3.5. Task-Specific Dataset Construction
@@
Table 4. Statistics of evaluation datasets for FinNI-eval and FinCL-eval. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
@@
3.5.1. FinNI-eval Dataset Construction
Following the annotation procedure in Section 3.3, we construct the FinNI-eval dataset with 28,787 instances derived from annotated sentences and tables. Each instance consists of an input block and an answer block. The input block includes a task instruction and the corresponding content, while the answer block is a JSON list of identified numerical values with their types (or an empty list).
@@ 3.5.1. FinNI-eval Dataset Construction
3.5.2. FinCL-eval Dataset Construction
Following the FinCL formulation in Section 3.1.2, we construct the FinCL-eval dataset with 261,457 query–answer pairs for numerical entity normalization, along with a US-GAAP taxonomy containing 17,688 concepts. Each query includes a numerical entity, its type, and surrounding context, while the answer is the corresponding US-GAAP concept.
@@ 3.6. Evaluation
We propose a unified evaluation framework for 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:. As shown in Figure 3, each input is decomposed into two subtasks, which together produce structured triplets. This design jointly evaluates zero-shot numerical extraction and concept alignment. For comparison with token-classification baselines, we report macro and micro Precision, Recall, and F1 (Sokolova and Lapalme, 2009). FinNI is evaluated using pair-level metrics, while FinCL is evaluated with accuracy.
FinNI Evaluation. FinNI uses pair-level metrics to assess whether models correctly extract numerical values from text-table inputs.
@@
FinCL Evaluation. FinCL is formulated as a retrieval–reranking task to avoid impractical extreme classification. We embed each taxonomy concept using text-embedding-3-small [^7] https://platform.openai.com/docs/models/text-embedding-3-small and retrieve the top-$h$ candidates from the US-GAAP taxonomy by semantic similarity. The LLM then reranks candidates using contextual reasoning to select the final tag.
Figure 3. A unified evaluation framework on 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark. Note: The Input content is drawn from the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: dataset, and the ”Shared LLMs” are used for identification in the FinNI task and reranking in the FinCL task.
3.6.1. Evaluation Models
@@
4. Experiment and Result
4.1. Overall Results
@@
Table 5. Overall Performance. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
Models Macro Micro
@@
GPT-4o 0.0865 0.0626 0.0569 0.1052 0.0731 0.0863
DeepSeek-V3 0.0949 0.0778 0.0659 0.1074 0.1264 0.1162
@@
Fino1-8B 0.0344 0.0143 0.0151 0.0419 0.0128 0.0197
BERT-large 0.0252 0.0266 0.0205 0.1518 0.1283 0.1391
@@
From a macro perspective, which emphasizes balanced performance across frequent and rare tags, large general-purpose LLMs clearly dominate. DeepSeek-V3, GPT-4o, and Llama-4-Scout achieve the strongest macro-F1 scores, surpassing all fine-tuned PLMs and indicating better generalization to long-tail concepts. Qwen3-32B also performs competitively, suggesting that model architecture and pretraining can compensate for a smaller scale. From a micro perspective, which reflects performance on frequent labels, the trend is similar: top LLMs remain competitive even without fine-tuning, while fine-tuned PLMs such as SECBERT and BERT-large excel mainly because they benefit from label frequency rather than broader coverage.
4.2. Subtask Results
Table 6 summarizes performance on the FinNI and FinCL subtasks. Overall, large LLMs outperform smaller models across both tasks in a zero-shot setting. On FinNI, DeepSeek-V3 achieves the best F1 score, driven by substantially higher recall, while GPT-4o performs competitively. In contrast, several models (e.g., Qwen3 variants) exhibit high precision but extremely low recall, indicating that they extract only a small fraction of target facts, which leads to poor F1 performance. Mid-sized models such as gemma-2-27B show more balanced behavior. In contrast, smaller models and the financial LLM Fino1-8B lag, suggesting that domain pretraining alone does not yield robust gains for numerical extraction.
@@ 4.2. Subtask Results
Table 6. Performance comparison of different models on the FinNI and FinCL subtasks. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
Model FinNI FinCL
Precision Recall F1 Accuracy
GPT-4o 0.5893 0.4977 0.5397 0.1829
DeepSeek-V3 0.5886 0.8430 0.6932 0.1889
Llama-4-Scout-17B-16E-Instruct 0.4668 0.3164 0.3771 0.1649
Llama-3.3-70B-Instruct 0.4826 0.3301 0.3920 0.1318
DeepSeek-R1-Distill-Qwen-32B 0.5676 0.1942 0.2894 0.1141
Qwen3-32B 0.6991 0.1804 0.2868 0.1277
gemma-2-27b-it 0.5060 0.4526 0.4778 0.1099
Qwen3-14B 0.6912 0.1487 0.2448 0.1144
Llama-3.1-8B-Instruct 0.3874 0.1761 0.2421 0.0913
Llama-3.2-3B-Instruct 0.1856 0.1203 0.1460 0.0415
Qwen3-1.7B 0.7362 0.0281 0.0541 0.0735
Qwen3-0.6B 0.2803 0.0019 0.0038 0.0414
Fino1-8B 0.3431 0.1293 0.1878 0.0807
FinCL remains significantly more challenging than FinNI. All FinCL results are reported under a retrieval–reranking setup, where models select the final tag from the top-200 retrieved US-GAAP candidates. Although retrieval results are not separately tabulated due to space constraints, structure-aware retrieval achieves over 0.65 Acc@200 for sentence-based entities but remains below 0.30 Acc@200 for table-based entities, resulting in an overall Acc@200 of approximately 0.33. This indicates that retrieval coverage is already limited, especially for tabular contexts, which imposes a ceiling on downstream reranking. Even under this constrained candidate space, the best-performing models achieve accuracy below 0.19, with DeepSeek-V3 and GPT-4o forming a narrow top tier. These results highlight a clear extraction–alignment gap: while modern LLMs can recover numerical facts with reasonable accuracy, fine-grained US-GAAP disambiguation remains intrinsically difficult, particularly for table-origin entities.
4.3. Ablation analysis